# Aligning Large Language Models to a Domain-specific Graph Database

### v2

@@ Abstract
Graph Databases (Graph DB) are widely applied in various fields, including finance, social networks, and medicine. However, translating Natural Language (NL) into the Graph Query Language (GQL), commonly known as NL2GQL, proves to be challenging due to its inherent complexity and specialized nature. Some approaches have sought to utilize Large Language Models (LLMs) to address analogous tasks like text2SQL. Nevertheless, when it comes to NL2GQL tasks on a particular domain, the absence of domain-specific NL-GQL data pairs makes it difficult to establish alignment between LLMs and the graph DB. To address this challenge, we propose a well-defined pipeline. Specifically, we utilize ChatGPT to create NL-GQL data pairs based on the given graph DB with self-instruct. Then, we use the created data to fine-tune LLMs, thereby achieving alignment between LLMs and the graph DB. Additionally, during inference, we propose a method that extracts relevant schema to the queried NL as the input context to guide LLMs for generating accurate GQLs. We evaluate our method on two constructed datasets deriving from graph DBs in finance domain and medicine domain, namely FinGQL and MediGQL. Experimental results demonstrate that our method significantly outperforms a set of baseline methods, with improvements of 5.90 and 6.36 absolute points on EM, and 6.00 and 7.09 absolute points on EX, respectively.
Aligning Large Language Models to a Domain-specific Graph Database
[…]
@@ 4.1 Generating NL-GQL Pairs as Instructions
To ensure the consistency of $Q^{\prime}$ and $Q^{{}^{\prime\prime}}$ across all data entries, we compare encode them via the sequence matching encoder and compare their similarities. If it falls below a certain threshold, we consider the $Q^{\prime}$ and $L^{\prime}$ to be inconsistent, leading to the discarding of the respective data entry. Otherwise, we include the NL-GQL pairs into data pool $Q^{\prime},L^{\prime}\in\mathcal{D}$.
@@ Appendix B Dataset Presentation
    MATCH (t:trade{name:"钢铁"(steel)})

### v3

@@ Abstract
Graph Databases (Graph DB) find extensive application across diverse domains such as finance, social networks, and medicine. Yet, the translation of Natural Language (NL) into the Graph Query Language (GQL), referred to as NL2GQL, poses significant challenges owing to its intricate and specialized nature. Some approaches have sought to utilize Large Language Models (LLMs) to address analogous tasks like text2SQL. Nonetheless, in the realm of NL2GQL tasks tailored to a particular domain, the absence of domain-specific NL-GQL data pairs adds complexity to aligning LLMs with the graph DB. To tackle this challenge, we present a well-defined pipeline. Initially, we utilize ChatGPT to generate NL-GQL data pairs, leveraging the provided graph DB with self-instruction. Subsequently, we employ the generated data to fine-tune LLMs, ensuring alignment between LLMs and the graph DB. Moreover, we find the importance of relevant schema in efficiently generating accurate GQLs. Thus, we introduce a method to extract relevant schema as the input context. We evaluate our method using two carefully constructed datasets derived from graph DBs in the finance and medicine domains, named FinGQL and MediGQL. Experimental results reveal that our approach significantly outperforms a set of baseline methods, with improvements of 5.90 and 6.36 absolute points on EM, and 6.00 and 7.09 absolute points on EX for FinGQL and MediGQL, respectively.
Aligning Large Language Models to a Domain-specific Graph Database for NL2GQL
Yuanyuan Liang1, Keren Tan1, Tingyu Xie2, Wenbiao Tao1 , Siyuan Wang1 Yunshi Lan1thanks: *Corresponding author,Weining Qian1 1 East China Normal University, 2 Zhejiang University {leonyuany,tankeren1020,wbtao,sywang}@stu.ecnu.edu.cn tingyuxie@zju.edu.cn, {yslan,wnqian}@dase.ecnu.edu.cn
@@ 4.1 Generating NL-GQL Pairs as Instructions
To ensure the consistency of $L$ and $Q$ across all data entries, we compare encode them via the sequence matching encoder and compare their similarities. If it falls below a certain threshold, we consider the $Q^{\prime}$ and $L^{\prime}$ to be inconsistent, leading to the discarding of the respective data entry. Otherwise, we include the NL-GQL pairs into data pool $Q^{\prime},L^{\prime}\in\mathcal{D}$.
@@ Appendix B Dataset Presentation
    MATCH (t:trade{name:"{CJK*}UTF8gbsn钢铁"(steel)})